# Machine-generated text detection prevents language model collapse

### v5

@@ 5 The impact of decoding strategies on model collapse
We carry out recursive training as described in section 4.1 on the open-ended text generation task by fine-tuning LLMs on the WikiText-2 dataset Merity et al. (2016) with GPT-2 $124$M  Radford et al. (2019) and SmolLM2 $360$Allal et al. (2025). The Wikipedia articles are segmented into non-overlapping chunks of $512$ tokens, where the first $256$ are used as the context ($\mathbf{x}$), and the remaining $256$ as the continuation ($\hat{\mathbf{x}}$). We conduct full fine-tuning for $1$ epoch and, to avoid information leakage between generation and training, define cross-entropy loss only on the generated sequence of each sample Dohmatob et al. (2024). Additional details can be found in Appendix A.3. We evaluate a range of decoding strategies to assess the effect on model collapse: greedy decoding ($\tau=0$), $5$-way beam search, pure sampling ($\tau=1$), temperature ($\tau=0.9$), top-$k$ (with $k=50$), and nucleus sampling ($\eta=0.95$). The hyperparameter settings for these decoding strategies were based on recommendations from prior work Holtzman et al. (2020); Shumailov et al. (2024); Arias et al. (2025).
Figure 1 depicts the perplexity and evaluation accuracy on the WikiText-2 test set for every model generation. Additionally, we obtain scores for the qualitative metrics using the outputs generated by the model (i.e. $\{\hat{\mathbf{x}}\}_{s=1}^{n}$ of $\mathcal{D}_{\text{S}}^{i}$ in Algorithm 1), and enumerate them in Table 1 for generations 00 and $9$. We observe that deterministic decoding leads to the most severe model collapse. While stochastic sampling methods exhibit linear degradation across generations, collapse accelerates under greedy decoding and beam search before plateauing in later generations. At generation 00, deterministic strategies yield significantly less fluent and more repetitive text, with MAUVE scores below $5\%$ and diversity scores less than $20\%$. The disparity in generation quality between deterministic and stochastic strategies in the open-ended text generation task has been explored in related literature Holtzman et al. (2020); Pillutla et al. (2021). Here, we demonstrate that this disparity compounds across recursive training, resulting in significantly higher perplexity at generation $9$. While deterministic methods are rarely used in open-ended generation, we included them for completeness and due to the choice of beam-search in the experiments of Shumailov et al. (2024), but exclude them in subsequent experiments due to their unrealistic collapse.
Pure sampling impacted the models differently. For GPT-2, sampling directly from the probability distribution produces diverse and fluent text at generation 00, but training recursively on these outputs results in the worst test perplexity among stochastic methods ($58.64$) and generated text that has low similarity to human text (MAUVE $7.18$). In contrast, with SmolLM2, pure sampling yields the smallest decline in task performance and maintains the closest overall similarity to human references across all evaluated metrics.
@@
In our subsequent experiments on preventing model collapse, we seek to validate that our method can work in the most extreme scenario. For this reason, we evaluate the models using the worst-performing stochastic decoding method (pure sampling for GPT-2 and temperature sampling for SmolLM2). In addition, to facilitate direct comparisons, we also evaluate with top-$k$ decoding due to the consistent performance across models.
@@ 6 Preventing model collapse
$\begin{split}\mathcal{D}^{i}\sim&\,\,{\text{sample}_{i\geq 1}}\,\,\mathcal{D}_{\text{H}},\alpha\\ &\,\,{\text{sample}_{i\geq 1}}\,\,\mathcal{D}_{\text{S}}^{i},\beta\\ &\,\,{\text{sample}_{i\geq 2}}\,\,\left\{\mathcal{D}_{\text{S}}^{i-1},\dots,\mathcal{D}_{\text{S}}^{1}\right\},\frac{\gamma}{(i-1)}\,,\end{split}$ (3)
@@
We explore the following settings: (i) fully synthetic $\left(\alpha\!=\!0,\beta\!=\!1,\gamma\!=\!0\right)$, where training data consists entirely of synthetic samples from the previous generation, (ii) partially synthetic $\left(\alpha\!>\!0,\beta\!=\!1,\gamma\!=\!0\right)$, where the same proportion of human data is added to the training data at every generation, and (iii) partially synthetic with synthetic data accumulated across generations $\left(\alpha\!=\!0.5,\beta\!=\!0.5,\gamma\!=\!0.5\right)$ as proposed in Gerstgrasser et al. (2024). We evaluate our method in the partially synthetic setting and vary the mixing coefficients $\alpha$, $\beta$, and $\gamma$, however, our method does not assume access to the values of the mixing coefficients and hence the data distribution. To prevent model collapse when the origin of each training sample is unknown, we train a machine-generated detector that estimates the likelihood of text origin (section 6.1). We then use this information to conduct importance sampling (section 6.2) that ultimately mitigates model collapse.
@@
Figure 2: Model collapse mitigation with GPT-2 or SmolLM2 under partially synthetic recursive training $\left(\alpha\!=\!1,\beta\!=\!1,\gamma\!=\!0\right)$ for generations 00 to $9$. The baseline is equivalent to training on all the data in the pool and the Oracle performance represents a perfect machine-generated text detector that filters all synthetic samples.
@@ 6.1 Machine-generated text detection performance
We trained and evaluated RoBERTa, DeBERTav3 and ModernBERT on the MAGE dataset Li et al. (2024), a machine-generated text detection benchmark based on documents from $10$ domains which have been used to generate text from $27$ LLMs. We adopt the preset training / validation / test splits ($80\%$/$10\%$/$10\%$). We also test on the more demanding out-of-distribution test set that contains human-curated text from $4$ unseen domains and machine-generated samples by an unseen LLM (GPT-4). Each model was fine-tuned for 5 epochs using a binary cross-entropy loss. More details on the model training can be found in Appendix B. Performance is enumerated in Table 2. ModernBERT [^3] Fine-tuned ModernBERT machine-generated text detector, huggingface.co/GeorgeDrayson/modernbert-ai-detection. yielded the best classification performance on both aforementioned test sets with an AUC of .986.986.986.986 and .943.943.943.943, respectively. This is comparable to the top-performing model evaluated by Li et al. (2024), Longformer, which achieved an in-distribution and out-of-distribution AUC of .99.99.99.99 and .94.94.94.94.
@@
RoBERTa .$982$940.940.940.940 .$940$$846$$806$804.804.804.804
DeBERTav3 .$971$$\mathbf{.954}$$\mathbf{.954}$$817$$812$810.810.810.810
ModernBERT .$\mathbf{.986}$948.948.948.948 .$948$$\mathbf{.943}$$\mathbf{.861}$$\mathbf{.860}$
@@
6.2 Informed sampling of training data
@@
Given a dataset at generation $i$, $\mathcal{D}_{i}\sim g(x)$, composed of an unknown mixture of human and synthetic samples, our goal is to sample a training dataset from a target human data distribution $\mathcal{D}_{\text{H}}\sim h(x)$ to prevent model collapse. We consider that a language model has collapsed if the inclusion of synthetic samples in the training data results in degraded performance compared to training exclusively on human samples.
@@
We use Sampling Importance Resampling (SIR) Rubin (1988), a method for approximately sampling from a target distribution $h(x)$ based on sampling with importance weights from a proposal distribution $g(x)$ using the normalised likelihood ratio $h(x)/g(x)$. As this ratio is intractable in our case, we instead employ a machine-generated text detector to assign each sample $\mathbf{x}_{i}$, $\forall i$, a predicted probability $q\!\left(\mathbf{x}_{i}\right)$ of being machine-generated, which we treat as an approximation for the likelihood ratio.
@@
As the detector has been trained on an unbalanced dataset ($29\%$ human samples), the predictions are biased towards attributing text as machine-generated, reflected in the optimal classification threshold of $0.8674$ (00/$1$: human/machine-generated). To ameliorate this, we apply a bias term $b\!\geq\!1$ (see Appendix C) to the probabilities, followed by normalising the weights using
$w_{i}=\frac{\left(1-q\!\left(\mathbf{x}_{i}\right)\right)^{b}}{\sum_{j=1}^{n}\left(1-q\!\left(\mathbf{x}_{j}\right)\right)^{b}}\,\,,$ (4)
@@
where $w_{i}\!\in\![0,1]$ denotes the weight for sample $\mathbf{x}_{i}$, $\forall i$. From the $n$ weighted training samples, we draw $k\!\times\!n$ samples with replacement, with $k=1.5$ to allow for a $50\%$ upsampling of the training data. In this way, we obtain a revised set of samples that we use in our recursive training regime.

### v6

@@ 5 The impact of decoding strategies on model collapse
We carry out recursive training as described in section 4.1 on the open-ended text generation task by fine-tuning GPT-2 124124M  Radford et al. (2019) and SmolLM2 360360Allal et al. (2025) on the WikiText-2 dataset Merity et al. (2017). The Wikipedia articles are segmented into non-overlapping chunks of $512$ tokens, where the first $256$ are used as the context ($\mathbf{x}$), and the remaining $256$ as the continuation ($\hat{\mathbf{x}}$). We conduct full fine-tuning for $1$ epoch and, to avoid information leakage between generation and training, define cross-entropy loss only on the generated sequence of each sample, $\mathcal{\hat{X}}\!=\!\{\mathbf{\hat{x}}_{1},\dots,\mathbf{\hat{x}}_{n}\}$ Dohmatob et al. (2024). Additional details can be found in Appendix A.3. We evaluate a range of decoding strategies to assess the effect on model collapse: greedy decoding ($\tau=0$), 55-way beam search, pure sampling ($\tau=1$), temperature ($\tau=0.9$), top-$k$ (with $k=50$), and nucleus sampling ($\eta=0.95$). The hyperparameter settings for these decoding strategies were based on recommendations from prior work Holtzman et al. (2020); Shumailov et al. (2024); Arias et al. (2025).
Figure 1 depicts the perplexity and evaluation accuracy on the WikiText-2 test set for every model generation. Additionally, we obtain scores for the qualitative metrics using the outputs generated by the model before being used for training (i.e. $\{\hat{\mathbf{x}}\}_{s=1}^{n}$ of $\mathcal{D}_{\text{S}}^{i}$ in Algorithm 1), and enumerate them in Table 1 for generations 0 and $9$.
We observe that deterministic decoding strategies lead to the most severe model collapse, with perplexity scores increasing by a factor of $\sim$$2.5$. Stochastic sampling methods exhibit linear degradation across generations, while collapse accelerates under greedy decoding and beam search before plateauing in later generations. Prior to recursive training (at generation 0), deterministic strategies yield significantly less fluent and more repetitive text, with MAUVE scores $5\%$ and diversity scores $20\%$. The disparity in data quality between deterministic and stochastic strategies in the open-ended text generation task has been explored in related literature Holtzman et al. (2020); Pillutla et al. (2021). Here, we demonstrate that this disparity compounds across recursive training, resulting in a significant drop in downstream performance and output quality at generation $9$. While deterministic methods are rarely used in open-ended generation, we included them to facilitate a better comparison with Shumailov et al. (2024) (where beam-search is used), but exclude them in subsequent experiments due to their unrealistic collapse.
@@
Figure 2: Recursive training process. The human data, $\mathcal{D}_{\text{H}}$, is used to train the first model iteration, $p^{0}$, which is then prompted to generate a synthetic dataset, $\mathcal{D}_{\text{S}}^{1}$. Subsequently, depending on the mixing coefficients $\alpha$, $\beta$, and $\gamma$, further model iterations are trained on a mix of human and synthetic data from the previous generation(s).
The effect of sampling directly from the probability distribution varied significantly between the two models. For GPT-2, pure sampling produces diverse and fluent text at generation 0, but training recursively on these outputs results in the worst test perplexity among stochastic methods ($58.64$), and generated text that has low similarity to human text (MAUVE of $7.18$). In contrast, with SmolLM2, pure sampling yields the smallest decline in task performance and maintains the closest overall similarity to human references across all evaluated metrics. Performance with top-$k$ sampling, on the other hand, was consistent across models. It led to the second-best task performance of all decoding strategies and, compared to the other truncation strategy, nucleus sampling, resulted in a smaller drop in diversity and stronger semantic resemblance to the human reference.
Temperature sampling led to the most repetitive outputs after recursive training, with diversity decreasing by $\sim$$70\%$ in both models. For SmolLM2, it also resulted in the greatest semantic divergence from human-generated text, indicating pronounced model collapse. While temperature is often deployed in open-ended text generation to improve data quality Holtzman et al. (2020), our results demonstrate that recursively training on synthetic data generated with temperature sampling can lead to model output convergence and a dramatic reduction in diversity.
In our subsequent experiments on preventing model collapse, we seek to validate that our method can work in the most challenging yet realistic scenario. For this reason, we evaluate the models using the worst-performing stochastic decoding method: pure sampling for GPT-2 and temperature sampling for SmolLM2. In addition, to facilitate direct comparisons, we also evaluate with top-$k$ decoding due to the consistent performance across models.
@@ 6 Preventing model collapse
[…]
@@
The recursive training process is depicted in Figure 2. We explore the following settings: (i) fully synthetic $\left(\alpha\!=\!0,\beta\!=\!1,\gamma\!=\!0\right)$, where training data consists entirely of synthetic samples from the previous generation, (ii) partially synthetic $\left(\alpha\!>\!0,\beta\!=\!1,\gamma\!=\!0\right)$, where the same proportion of human data is added to the training data at every generation, and (iii) partially synthetic with synthetic data accumulated across generations $\left(\alpha\!>\!0,\beta\!>\!0,\gamma\!>\!0\right)$ as proposed in Gerstgrasser et al. (2024). We evaluate our method in the partially synthetic setting and vary the mixing coefficients $\alpha$, $\beta$, and $\gamma$. However, our method does not assume access to the values of the mixing coefficients, and hence the data distribution. To prevent model collapse when the origin of each training sample is unknown, we train a machine-generated text detector. The detector estimates the likelihood that a text sample is generated by a human (section 6.1). We then use this information to conduct importance resampling on the training data (section 6.2) that ultimately mitigates model collapse.
@@
Figure 3: Model collapse mitigation with GPT-2 or SmolLM2 under partially synthetic recursive training $\left(\alpha\!=\!1,\beta\!=\!1,\gamma\!=\!0\right)$ for generations 0 to $9$. The baseline is equivalent to training on all the data in the pool and the ‘Oracle’ performance represents a perfect machine-generated text detector that filters all synthetic samples.
@@ 6.1 Machine-generated text detection performance
For the machine-generated text detector, we trained and evaluated three pretrained models (RoBERTa, DeBERTav3, and ModernBERT) on the MAGE dataset Li et al. (2024), a machine-generated text detection benchmark based on documents from $10$ domains. For each human-written text sample in the dataset, $27$ LLMs were prompted with the first $30$ words to generate a set of machine-generated samples. We adopt the preset training / validation / test splits ($80\%$$\%/1010\%$). We also test on the more demanding out-of-distribution test set that contains human-curated text from $4$ unseen domains and machine-generated samples by an unseen LLM (GPT-4). Each model is fine-tuned for $5$ epochs using a binary cross-entropy loss and the best checkpoint is selected based on validation performance. More details on the model training can be found in Appendix B. Performance is enumerated in Table 2. ModernBERT [^4] Our fine-tuned ModernBERT machine-generated text detector is available at huggingface.co/GeorgeDrayson/modernbert-ai-detection. yielded the best classification performance on both test sets with an AUC of .986.986 and .943.943, respectively. This is comparable to the top-performing model evaluated by Li et al. (2024), Longformer, which achieved an in-distribution and out-of-distribution AUC of .99.99 and .94.94, respectively.
@@
RoBERTa .982.982 .940.940 .940.940 .846.846 .806.806 .804.804
DeBERTav3 .971.971 $\mathbf{.954}$ $\mathbf{.954}$ .817.817 .812.812 .810.810
ModernBERT $\mathbf{.986}$ .948.948 .948.948 $\mathbf{.943}$ $\mathbf{.861}$ $\mathbf{.860}$
@@
6.2 Informed resampling of training data
@@
Given a dataset at generation $i$, $\mathcal{D}_{i}\sim g(x)$, composed of an unknown mixture of human and synthetic samples, our goal is to sample a training dataset from a target human data distribution $\mathcal{D}_{\text{H}}\sim h(x)$ to prevent model collapse. We consider that a language model has collapsed when the inclusion of synthetic samples in the training data results in degraded performance compared to training exclusively on human samples.
[…]
GPT-2 top-$k$ .5.5, 11, 0 $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-7.28\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+2.37\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+4.54\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-1.77\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+0.46\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+9.03\%$}
.5.5, .5.5, .5.5 $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-5.94\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+1.72\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+2.43\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-4.28\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+3.99\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+10.16\%$}
11, 11, 0 $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-4.45\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+1.49\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+3.59\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-3.58\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+1.36\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+6.76\%$}
pure sampling .5.5, 11, 0 $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-7.41\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+1.50\%$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-1.30\%$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}+36.71\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+74.06\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+50.23\%$}
.5.5, .5.5, .5.5 $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-6.54\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+1.50\%$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-1.07\%$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}+21.39\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+16.38\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+25.45\%$}
11, 11, 0 $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-25.65\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+0.96\%$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-0.71\%$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}+20.70\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+16.42\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+20.99\%$}
SmolLM2 360360M top-$k$ .5.5, 11, 0 $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-4.60\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+0.68\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+7.06\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-1.17\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+15.28\%$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-0.73\%$}
.5.5, .5.5, .5.5 $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-4.37\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+0.62\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+4.25\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-0.42\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+2.05\%$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-0.74\%$}
11, 11, 0 $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-3.05\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+0.54\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+6.20\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-1.74\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+10.62\%$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-0.61\%$}
temperature .5.5, 11, 0 $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-3.91\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+0.55\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+14.62\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-1.46\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+20.92\%$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-1.86\%$}
.5.5, .5.5, .5.5 $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-3.24\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+0.48\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+7.44\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-1.98\%$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-0.33\%$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-1.35\%$}
11, 11, 0 $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-2.23\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+0.39\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+8.55\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-1.80\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+13.98\%$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-0.34\%$}
SmolLM2 135135M top-$k$ 11, 11, 0 $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-4.19\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+0.75\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+9.96\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-2.15\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+13.08\%$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-0.76\%$}
SmolLM2 1.71.7B top-$k$ 11, 11, 0 $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-0.85\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+0.06\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+3.44\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-1.11\%$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+8.46\%$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-0.65\%$}
Table 3: Percentage of change in data quality and model performance when using our proposed mitigation strategy versus the baseline (final model generation). Results are shown for top-$k$ decoding and pure sampling/temperature for different values of $\alpha$, $\beta$, and $\gamma$ (blue / red: positive / negative results, $\uparrow$ / $\downarrow$: higher / lower is better).
@@
We use Sampling Importance Resampling (SIR) Rubin (1988), a method for approximately sampling from a target distribution $h(x)$ based on sampling with importance weights from a proposal distribution $g(x)$ using the normalised likelihood ratio $h(x)/g(x)$. As this ratio is intractable in our case, we instead employ a machine-generated text detector to assign each sample, $\mathbf{x}_{j}$, a predicted probability $q\!\left(\mathbf{x}_{j}\right)$ of being machine-generated, which we treat as an approximation for the likelihood ratio.
@@
As the detector has been trained on an unbalanced dataset ($29\%$ human samples), the predictions are biased towards attributing text as machine-generated, reflected in the optimal classification threshold of $0.8674$ (0/11: human/machine-generated). To ameliorate this, we apply a bias term $b\!\geq\!1$ (see Appendix C) to the probabilities, followed by normalising the weights using
$w_{j}=\frac{\left(1-q\!\left(\mathbf{x}_{j}\right)\right)^{b}}{\sum_{l=1}^{n}\left(1-q\!\left(\mathbf{x}_{l}\right)\right)^{b}}\,\,$, (4)
@@
where $w_{j}\!\in\![0,1]$ denotes the weight for sample $\mathbf{x}_{j}$, $\forall j$. From the $n$ weighted training samples, we draw $k\!\times\!n$ samples with replacement, with $k=1.5$ to allow for a $50\%$ upsampling of the training data. In this way, we obtain a revised set of samples that we use in our recursive training regime.